# TEAL: New Selection Strategy for Small Buffers in Experience Replay Class Incremenal Learning

### v1

@@ 4 Empirical evaluation
We report two settings: (i) Integrated:we evaluate the beneficial contribution of TEAL to competitive ER-IL methods, by replacing their native selection strategy with TEAL (Section 4.2.1); (ii) Alternative selection strategies:we evaluate the beneficial contribution of TEAL as compared to Herding (Section 4.2.2). In both cases, we incrementally train a deep-learning model with a fixed-size replay buffer and monitor the average accuracy defined below upon completion of each task $t$.
@@ 4.1 Methodology
The majority of the experiments are conducted using the open-source CL library Avalanche [16], with the exception of those involving XDER [1], which is not integrated into Avalanche. For XDER, we utilize the code provided by the authors. In order to guarantee a fair comparison and in all experiments, we employ identical network architectures and maintain consistent experimental conditions.
@@
We make certain in both settings that the buffer remains class-balanced, namely, each update maintains an equal representation of exemplars across all classes. When the buffer size is not evenly divisible by the number of classes, there may be instances where certain classes possess one additional exemplar. To maintain this balance, we follow this procedure: Prior to adding new exemplars we calculate $n=\frac{|{\mathcal{M}}|}{N^{t}}$, where $N^{t}$ is the total number of existing and new classes. Subsequently, we adjust the number of exemplars from existing classes to accommodate $n$ by removing the least likely points, followed by the addition of $n$ exemplars from each new class.
@@
In the second setting, we employ a simple baseline ER-IL model. This model updates a fixed-size buffer of exemplars after training each task $t$ and replays it while training on task $t+1$. Additionally, we utilize a weighted data loader to ensure a balanced mix of data from both the new classes and the exemplars stored in the buffer in each batch. Across experiments, the only variation lies in the selection strategy employed: some experiments utilize random-sampling, others utilize Herding, and the remaining experiments leverage TEAL. In some experiments, we replace the native buffer selection strategy of an existing ER-IL method with TEAL or Herding to assess their impact. Other than this change in the method’s buffer population mechanism, everything else remains the same.
Datasets    We use several well known Continual Learning benchmarks. (i) Split CIFAR-100[22, 5], a dataset created by splitting CIFAR-100 [13] into 10 tasks, each containing 10 different classes of 32X32 images, with 500 images per class for training and 100 for testing. (ii) Split tinyImageNet, a dataset created by splitting tinyImageNet [14] into 10 tasks, each containing 20 different classes of 64X64 images, with 500 images per class for training and 50 for testing. (iii) Split CUB-200[3], a dataset created by splitting the CUB-200 [26] high-resolution image classification dataset, consisting of 200 categories of birds, with around 30 images per class for training and 30 for testing.
Architectures    In the first setting, except for one experiment, we employ a ResNet-18 model in all our experiments. The exception is the experiment involving the training of XDER on the Split CUB-200 dataset. Due to computational limitations, we used instead a pre-trained ResNet-50 model [11], and this was maintained under all the relevant conditions. In the second setting, we use a smaller version of ResNet-18 [11] as a simple baseline ER-IL model [see 17].
Metrics    As customary, we employ a score that reflects the accuracy during each stage of the incremental training. Let $T$ denote the total number of tasks, and $a_{t,i}$ denote the accuracy of task $i$ at the end of task $t$ ( $i\leq t\leq T$). For every task $t={1,\dots,T}$, define it average accuracy $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i}$. This metric provides a single value for each incremental step, enabling us to directly compare different methods at each step. Since all of the experiments we conduct are task-balanced in the sense that each task has an equal number of classes, there is no need to add a weighting term to the metric.
Baselines    The following methods are suitable for our investigation as they are designed to address ER-IL: XDER [1] updates the memory buffer by adding current information to past memories and preparing the model for new tasks. ER-ACE [2], or ER with Asymmetric Cross-Entropy, uses a separate loss for new and past tasks. BiC [28] introduces a bias correction layer to address class imbalance in incremental learning. iCaRL [22] combines the incremental classifier and representation learning with a nearest-mean-of-exemplars strategy, and Herding for exemplar selection. Since the classifier is based on the nearest-mean-of-exemplars, Herding is crucial for the selection strategy, as it guarantees that the features of the exemplars in the buffer approximate the original classes means. GEM [17] uses Gradient Episodic Memory to prevent gradient leakage between new and past tasks through constrained optimization. GDumb [21] employs a greedy sampling strategy for memory storage and retrains the neural network from scratch during inference. ER [4] explores different selection strategies in a simple ER framework.
In order to evaluate the competitiveness of these methods, we conducted multiple runs on Split CIFAR-100, utilizing fixed buffer sizes of 300, 500, and 2000. Fig. 2 shows the results.
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Probably because the memory buffer is so small, and as can be seen in Fig. 2, GEM and GDumb perform poorly in a relative comparison. BiC performs well on buffer size 2000, but significantly degrades for smaller buffers for which it is not suited. iCarl demonstrates high performance, but it is highly non-modular, being heavily dependent on the Herding selection strategy while using a K-Means classifier. Hence we cannot integrate it with TEAL.
@@
We are therefore left with methods XDER, ER-ACE and ER, which are both suitable and competitive, to be examined with and without TEAL as the selection mechanism. For ER-ACE, since it depends on populating the buffer with exemplars at the beginning of the task before training on the new classes, we cannot rely on the representation provided by a trained model. Therefore, we adjust the integration of TEAL by filling the buffer in two stages: First, at the beginning of task $t$, we update the buffer with exemplars from the new classes using the random-sampling selection strategy. Second, after training task $t$, we replace the exemplars from the new classes in the buffer with exemplars chosen using TEAL. It’s important to note that throughout this process, the buffer maintains the fixed size of exemplars it allowed, ensuring compliance with the class-incremental framework.
@@ 4.2 Main results
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
Split CIFAR-100
(a) $|{\mathcal{M}}|=300$
(e) $|{\mathcal{M}}|=1000$
(f) $|{\mathcal{M}}|=2000$
Split tinyImageNet
(d) $|{\mathcal{M}}|=400$
(h) ER
(i) XDER
Split CUB-200
(g) ER-ACE
Figure 3: The performance of ER-IL methods with and without TEAL. Each plot displays the average accuracy after training incrementally on a different number of classes. Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy. The error bars correspond to standard error based on 4-10 repetitions.
4.2.1 TEAL integrated with SOTA ER-IL methods
Split CIFAR-100
(a) $|{\mathcal{M}}|=300$
(b) $|{\mathcal{M}}|=500$
(c) $|{\mathcal{M}}|=2000$
(e) $|{\mathcal{M}}|=1000$
(f) $|{\mathcal{M}}|=2000$
Split tinyImageNet
(d) $|{\mathcal{M}}|=400$
(h) ER
(i) XDER
Split CUB-200
(g) ER-ACE
Figure 4: Performance improvement of TEAL when integrated with different ER-IL methods. Each graph depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks. The first two rows present the three methods improvement on Split CIFAR-100 and Split tinyImageNet with different fixed buffer sizes denoted as $|{\mathcal{M}}|$. The last row pertains to Split CUB-200, where each graph showcases different buffer sizes for a different method.
Table 1: Final average accuracy for method with and without TEAL
Dataset Split CIFAR-100 Split tinyImageNet Split CUB-200
$|{\mathcal{M}}|$ 300 500 2000 400 1000 2000 200 400
XDER 41.97±0.25 47.97±0.22 57.69±0.21 20.49±0.19 32.68±0.24 39.76±0.31 43.63±1.78
XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17 23.61±0.32 34.28±0.24 40.76±0.23 49.55±0.47
Improvement 3.08 2.32 1.1 3.13 1.59 1.0 5.96
ER-ACE 35.63±0.36 40.6±0.34 53.6±0.17 13.02±0.2 17.37±0.34 21.71(±0.17) 8.56±0.15 11.25±0.99
ER-ACE+TEAL 40.59±0.2 44.25±0.35 56.11±0.19 14.43±0.29 18.88±0.21 24.12±0.24 10.41±0.16 12.33±0.26
Improvement 4.97 3.65 2.51 1.41 1.51 2.4 1.85 1.08
ER 14.82±0.21 18.21±0.29 36.54±0.31 7.75±0.06 7.7±0.06 8.27±0.08 8.98±0.18 12.01±0.52
ER+TEAL 17.97±0.25 22.39±0.31 38.17±0.37 7.86±0.05 7.99±0.06 8.82±0.09 10.37±0.23 14.13.76±0.3
Improvement 3.15 4.18 1.62 0.11 0.29 0.56 1.39 2.12
As mentioned above, we investigate the baseline methods XDER, ER-ACE and ER, comparing their performance when using either their native selection strategies or TEAL. We examined Split CIFAR-100 with buffer sizes of 300, 500 and 2000, Split tinyImageNet with buffer sizes of 400, 1000 and 2000, and Split CUB-200 with buffer sizes of 200222Due to computational limitation of the relevant package, we did not run XDER with buffer size 200. and 400. Fig. 3 shows the average accuracy $A_{t}$ after each task $t$. Fig. 4 shows the corresponding improvement of TEAL. Finally, the results of the final average accuracy $A_{t}$ and the differences ’improved-vanilla’ are reported in Table 1.
@@ 4.2.1 TEAL integrated with SOTA ER-IL methods
Table 1 shows that the difference (’improvement’ row) is always positive, namely, TEAL always improves the outcome. Looking at the improvement row for the XDER method, it is evident that the improvement provided by TEAL increases as the buffer size decreases. This trend is also demonstrated for ER-ACE and partially for ER on Split CIFAR-100.
4.2.2 Comparison of selection strategy
We begin with a simple ER-IL model on Split CIFAR-100333The order of the classes and the partition to tasks is randomly selected and fixed for all comparisons. with various buffer sizes, and three different selection strategies: random, Herding and TEAL. As depicted in Fig. 5, our method enhances performance as compared to random selection by almost 3% for smaller buffers (300, 400, 500) and about 2.5% for larger buffers (1000, 2000). When compared to Herding, our method is more effective for smaller buffer sizes, enhancing performance by up to 1.2% for a buffer size of 300, but less effective with large buffers. Similar results, using different class ordering, are shown in App A.1.
Figure 5: Split CIFAR-100: performance difference between TEAL and either random-sampling or Herding selection. Each bar corresponds to a fixed buffer size.
Dataset Split CIFAR-100
$|{\mathcal{M}}|$ 300 500 2000
XDER+TEAL_OneTime 44.13±0.26 49.87±0.22 59.05±0.14
XDER+TEAL 45.05±0.24 50.29±0.2 58.79±0.17
ER-ACE+TEAL_OneTime 35.81±0.42 40.16±0.3 51.52±0.24
ER-ACE+TEAL 36.4±0.12 40.26±0.23 51.68±0.35
ER+TEAL_OneTime 17.64±0.12 22.01±0.18 40.65±0.32
ER+TEAL 18.1±0.1 22.07±0.13 39.96±0.23
Table 2: The improvement of 2 variants of TEAL when integrated with different CIL methods: XDER, ER-ACE and ER. In almost all cases, the original TEAL matches or outperforms the OneTime variant.
Given the results shown in Fig. 5, wherein TEAL seems less effective than herding with large buffers when training the simple baseline model, we repeat this comparison while integrating both methods - TEAL and Herding - into competitive ER-IL methods. Fig. 6 shows the difference between the improvement obtained by TEAL and the one obtained by Herding. Since in almost all cases this difference is positive, we may conclude that TEAL is significantly more beneficial than Herding.

### v2

@@ 4 Empirical evaluation
We report two settings: (i) Integrated:we evaluate the beneficial contribution of TEAL to competitive ER-IL methods, by replacing their native selection strategy (vanilla version) with TEAL (Section 4.2); (ii) Alternative selection strategies:we evaluate the beneficial contribution of TEAL as compared to other selection strategies (Section 4.2). In both cases, we incrementally train a deep model with a fixed-size replay buffer and monitor the average accuracy defined below upon completion of each task $t$.
@@ 4.1 Methodology
The majority of the experiments are conducted using the open-source Continual Learning library Avalanche (Lomonaco et al. 2021), with the exception of those involving XDER (Boschini et al. 2022), which is not integrated into Avalanche. For XDER, we utilize the code provided by the authors. In order to guarantee a fair comparison and in all experiments, we employ identical network architectures and maintain consistent experimental conditions.
@@
We make certain in both settings that the buffer remains class-balanced, namely, each update maintains an equal representation of exemplars across all classes [^2] When the buffer size is not evenly divisible by the number of classes, there may classes with an additional exemplar.. To maintain balance, we follow this procedure: Prior to adding new exemplars we calculate $n=\frac{|{\mathcal{M}}|}{N^{t}}$, where $N^{t}$ is the total number of existing and new classes. Subsequently, we adjust the number of exemplars from existing classes to accommodate $n$ by removing redundant points, followed by the addition of $n$ exemplars from each new class.
@@
In the second setting, we employ a simple baseline ER-IL model. This model updates a fixed-size buffer of exemplars after training each task $t$ and replays it while training on task $t+1$. Additionally, we utilize a weighted data loader to ensure a balanced mix of data from both the new classes and the exemplars stored in the buffer in each batch. Across experiments, the only variation lies in the selection strategy employed: some experiments utilize one of the selection strategies baselines described below, and the remaining experiments employ TEAL.
Other than this change in the method’s buffer population mechanism, everything else remains the same.
ER-IL baselines    The following ER-IL methods are used: XDER (Boschini et al. 2022), which updates the memory buffer by integrating current and past information; ER-ACE (Caccia et al. 2021), which applies separate losses for new and past tasks; ER (Chaudhry et al. 2019a), which explores selection strategies in a simple ER framework; BiC (Wu et al. 2019); iCaRL (Rebuffi et al. 2017); GEM (Lopez-Paz and Ranzato 2017); and GDumb (Prabhu, Torr, and Dokania 2020).
To assess the suitability of these methods under the CIL conditions studied here, we used Split CIFAR-100 and set the buffer size to 300, 500, and 2000. Figure 3 presents the results for a buffer size of 300 (see Suppl B for results with larger buffer sizes). As clearly illustrated in Fig. 3, GEM and GDumb perform poorly in this scenario, likely due to their unsuitability for a very small memory buffer. Similarly, BiC remains competitive only when the buffer size is 2000. While iCarl remains competitive across all buffer sizes, it is inherently non-modular, relying heavily on its native Herding selection strategy and a K-Means classifier. Hence we cannot integrate it with TEAL.
@@
We are therefore left with the methods XDER, ER-ACE, and ER, which are both suitable and competitive for further examination with and without TEAL as the selection mechanism. For ER-ACE, which requires populating the buffer with exemplars at the start of the task before training on new classes, we cannot rely on the representation provided by a trained model. To address this, we adjust the integration of TEAL by filling the buffer in two stages. First, at the beginning of task $t$, we populate the buffer with exemplars from the new classes using a random-sampling selection strategy. Second, after completing the training for task $t$, we replace the exemplars from the new classes in the buffer with those selected using TEAL. It’s important to note that throughout this process, the buffer maintains a fixed size, ensuring compliance with the class-incremental framework.
Figure 3: Baseline CIL using Split CIFAR-100 with a small buffer size $|{\mathcal{M}}|=300$. 7 baseline methods are shown, reporting the average accuracy $A_{t}$ as new tasks are learned.
Selection Strategies baselines    We compare with the following selection strategies (see discussion of related work in the introduction): Random sampling, Herding, Uncertainty, and Centered (closest-to-center) - another common strategy that selects the exemplars closest to the center of all elements in feature space. Centered focuses on individual exemplars, unlike Herding, which keeps the mean of the whole constructed set close to the center.
Metrics    As customary, we employ a score that reflects the accuracy during each stage of the incremental training. Let $T$ denote the total number of tasks, and $a_{t,i}$ denote the accuracy of task $i$ at the end of task $t$ ( $i\leq t\leq T$). For every task $t={1,\dots,T}$, define its average accuracy $A_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i}$. This metric provides a single value for each incremental step, enabling us to directly compare different methods at each step. Since by construction each task has the same number of classes, there is no need for additional weighting terms. We do not include the metric of forgetting, which estimates how much the model has forgotten about previous tasks, since it is less suitable for the class-incremental setting: with the addition of new classes, performance inevitably drops across all classes (Masana et al. 2022).
Datasets    We use several well known Continual Learning benchmarks. (i) Split CIFAR-100(Rebuffi et al. 2017; Chaudhry et al. 2019b), a dataset created by splitting CIFAR-100 (Krizhevsky, Hinton et al. 2009) into 10 tasks, each containing 10 different classes of 32X32 images, with 500 images per class for training and 100 for testing. We also use a variation of this dataset by splitting CIFAR-100 into 20 tasks instead of 10, as reported in Suppl B. (ii) Split tinyImageNet, a dataset created by splitting tinyImageNet (Le and Yang 2015) into 10 tasks, each containing 20 different classes of 64X64 images, with 500 images per class for training and 50 for testing. (iii) Split CUB-200(Chaudhry et al. 2018b), a dataset created by splitting into 20 tasks the CUB-200 (Wah et al. 2011) high-resolution image classification dataset, consisting of 200 categories of birds, with around 30 images per class for training and 30 for testing.
Architectures    In the first setting, except for one experiment, we employ a ResNet-18 model in all our experiments. The exception is the experiment involving the training of XDER on the Split CUB-200 dataset. Due to computational limitations, we used instead a pre-trained ResNet-50 model (He et al. 2016), and this was maintained under all the relevant conditions. Additionally, we run some experiments with another architecture called ArchCraft, a ResNet variant designed for improved performance in CL (Lu et al. 2024). In the second setting, we use a smaller version of ResNet-18 (He et al. 2016) as a simple baseline ER-IL model (see Lopez-Paz and Ranzato 2017).
@@ 4.2 Main results
(a) Split CIFAR-100, $|{\mathcal{M}}|=300$
(b) Split tinyImageNet, $|{\mathcal{M}}|=400$
(c) Split CUB-200
Figure 4: The performance of ER-IL methods with and without TEAL. First row displays the average accuracy after training incrementally on a different number of classes. Each color corresponds to a different ER-IL method, where the continuous line represents the vanilla method, while the dashed line represents the method with TEAL as its selection strategy. The error bars correspond to standard error based on 4-10 repetitions. Second row depicts the difference in accuracy between TEAL and another method (XDER, ER-ACE, and ER) across all tasks.
TEAL integrated with SOTA ER-IL methods
Table 1: Final average accuracy for method with and without TEAL.
XDER ER-ACE ER
Dataset $|{\mathcal{M}}|$ Vanilla +TEAL Improvement Vanilla +TEAL Improvement Vanilla +TEAL Improvement
Split CIFAR-100 100 27.98±0.18 31.08±0.29 3.1 24.37±0.42 31.06±0.27 6.9 10.31±0.03 10.24±0.03 -0.07
300 41.97±0.25 45.05±0.24 3.08 35.37±0.29 41.28±0.2 5.92 14.82±0.21 17.97±0.25 3.15
500 47.97±0.22 50.29±0.2 2.32 40.7±0.25 45.99±0.33 5.29 18.21±0.29 22.39±0.31 4.18
1000 53.69±0.32 55.02±0.34 1.33 48.16±0.14 51.85±0.22 3.69 26.66±0.81 28.58±0.65 1.92
2000 57.69±0.21 58.79±0.17 1.1 55.99±0.1 58.7±0.17 2.71 36.54±0.31 38.17±0.37 1.62
3000 58.96±0.25 60.09±0.25 1.13 59.75±0.25 62.3±0.19 2.55 43.72±1.63 46.03±1.49 2.32
4000 59.89±0.16 60.42±0.28 0.54 62.6±0.22 64.81±0.2 2.22 49.03±0.89 52.86±0.62 3.82
Split tinyImageNet 200 14.6±0.16 16.76±0.3 2.16 11.3±0.08 12.47±0.22 1.17 7.95±0.02 8.01±0.03 0.07
400 20.42±0.12 23.58±0.18 3.16 13.02±0.2 14.43±0.29 1.41 7.75±0.06 7.86±0.05 0.11
600 25.74±0.09 28.71±0.23 2.97 14.48±0.13 15.7±0.14 1.22 7.54±0.02 7.8±0.07 0.26
1000 32.68±0.24 34.28±0.24 1.59 17.37±0.34 18.88±0.21 1.51 7.7±0.06 7.99±0.06 0.29
2000 39.76±0.31 40.76±0.23 1.0 21.71±0.17 24.12±0.24 2.4 8.27±0.08 8.82±0.09 0.56
4000 43.91±0.14 44.42±0.26 0.51 27.15±0.13 29.62±0.18 2.47 11.61±0.1 13.59±0.21 1.99
6000 44.71±0.15 45.44±0.39 0.73 30.3±0.23 33.3±0.23 3 16.53±0.15 19.07±0.18 2.54
Split CUB-200 200 8.56±0.15 10.41±0.16 1.85 8.98±0.18 10.37±0.23 1.39
400 43.63±1.78 49.55±0.47 5.96 11.25±0.99 12.33±0.26 1.08 12.01±0.52 14.13±0.3 2.12
As mentioned above, we investigate the baseline methods XDER, ER-ACE, and ER, comparing their performance using either their native selection strategies or TEAL. We conducted experiments on Split CIFAR-100 with buffer sizes ranging from 100 to 4000, Split tinyImageNet with buffer sizes from 200 to 6000, and Split CUB-200 with buffer sizes of 200333Due to computational limitations of the relevant package, we did not run XDER with a buffer size of 200. and 400. The results for final average accuracy $A_{t}$ and the differences between ’improved’ and ’vanilla’ are presented in Table 1. First row of Fig. 4 shows the average accuracy $A_{t}$ after each task $t$, while the second row illustrates the corresponding improvement from TEAL in selected experiments.
@@
Table 1 shows that the difference (indicated in the ’improvement’ row) is consistently positive, demonstrating that TEAL always enhances performance. However, it is important to note that when catastrophic forgetting in the vanilla version is too severe, the integration of TEAL can no longer mend the damage. Consequently, the enhancements to ER and ER-ACE on the Split tinyImageNet dataset are mostly notable with larger buffer sizes.
Note that the improvement provided by TEAL tends to increase as the buffer size decreases. To further explore the relationship between buffer size and relative improvement, we present a graph showing the final average accuracy improvement across various buffer sizes. The results for Split CIFAR-100 and Split tinyImageNet using XDER are illustrated in Fig. 5 (see results with ER-ACE and ER in Suppl B). Clearly, while improvements do occur across all buffer sizes, larger improvements are seen for smaller buffers.
Figure 5: Performance improvement of TEAL when integrated with XDER over various buffer sizes.
Comparison of selection strategy
We begin with a simple ER-IL model on Split CIFAR-100444The order of the classes and the partition to tasks is randomly selected and fixed in all conditions. with various buffer sizes, and 5 different selection strategies: random, Herding, Uncertainty, Centered and TEAL. As depicted in Fig. 7, our method enhances performance as compared to Uncertainty by 3-4%, and random selection by almost 3% for smaller buffers (300, 400, 500) and about 2.5% for larger buffers (1000, 2000). When compared to Herding, our method is more effective with smaller buffer sizes, enhancing performance by up to 1.2% for a buffer size of 300, but less effective with large buffers. It is possible that the mechanism of retaining a set of the nearest neighbors around the average sample in each class becomes less effective when the buffer is too small. Additionally, Herding lacks a component for selecting a diverse set, which may lead to the retention of similar exemplars, an issue that becomes more pronounced with smaller buffer sizes. Interestingly, despite its similarity to Herding, Centered performs worse on larger buffer sizes, resulting in a 1.5-2.5% improvement by TEAL. Similar results, using different class ordering, are shown in Suppl A.2.
@@
Figure 7: Split CIFAR-100: TEAL performance gain compared to 4 baselines: random-sampling, Herding, Centered, and Uncertainty, each bar corresponds to a fixed buffer size. The error bars correspond to standard error based on 10-20 repetitions.
Given the results shown in Fig. 7, we repeat this comparison while integrating both methods - TEAL and Herding - into competitive ER-IL methods. Fig. 6 shows the difference between the improvement obtained by TEAL and the one obtained by Herding. Since in almost all cases this difference is positive, we may conclude that TEAL is significantly more beneficial than Herding.
Results of Task-IL
The incorporation of TEAL into existing ER-IL method is also beneficial in the task incremental scenario, although the advantage is less pronounced, likely because the scenario is easier. Results are shown in Table 2.
Table 2: Task-IL, accuracy with and without TEAL
Dataset $|{\mathcal{M}}|$ XDER XDER+TEAL Improvement
Split CIFAR-100 100 73.75±0.59 76.56±0.27 2.81
300 83.5±0.06 84.46±0.26 0.96
500 85.96±0.14 86.1±0.2 0.14
Split tinyImageNet 200 42.48±0.38 46.55±0.37 4.07
400 55.03±0.15 59.2±0.47 4.17
600 63.69±0.21 65.26±0.28 1.57
1000 69.72±0.31 70.61±0.23 0.89